OpenStudy (anonymous):

solve for A, show work 18 - 6a = 4a - 4a + 12

OpenStudy (anonymous):

so add up the values of a again, just pretend a is the same situation as x. Just a different letter right?

OpenStudy (anonymous):

add 6a to the right, or the ones from the right to the left?

OpenStudy (anonymous):

doesn't matter which way, just pick which side you want to go for. So if I went with the left side. I would keep -6a, subtract 4a, and add 4a.

OpenStudy (anonymous):

if I did the right side, it would be the opposite; make sense?

OpenStudy (anonymous):

example with the problem if you would?

OpenStudy (anonymous):

So; you have: \[18 - 6a = 4a - 4a + 12\] if you wanted to go to the left hand side you can eliminate the variables on the right like so: \[18 - 6a - 4a + 4a = 4a - 4a - 4a + 4a + 12\] which can be simplified to \[18 - 6a = 12\]

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

so many 4a o.o

OpenStudy (anonymous):

Yes; however, this problem has a shortcut. If you notice \[18 - 6a = 4a - 4a + 12\] the \[4a - 4a + 12\] can be simplified just as \[12\]

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

yes ^.^

OpenStudy (anonymous):

18 - 6a + 4a = 12

OpenStudy (anonymous):

So since you can just simplify it right away, you an rewrite it without those because they cancel eachother out. so you start with \[18 - 6a = 12\] which is what we got before.

OpenStudy (anonymous):

Where'd you get the extra 4a?

OpenStudy (anonymous):

18 - 6a = 4a - 4a + 12 18 - 6a + 4a = 12 i just brought it over

OpenStudy (anonymous):

No, haha, I wasn't clear.

OpenStudy (anonymous):

Think of it this way. \[18 - 6a = 4a - 4a + 12\] is simple because the 4a and the -4a negate each other and cancel out because they are exact opposites. make sense?

OpenStudy (anonymous):

So since they are exact opposites; you can literally just pretend they don't even exist. It's like saying add one and subtract one. you're left with nothing.

OpenStudy (anonymous):

ಠ_ಠ so 18 - 6a = 12 the 4a and -4a just dont exsist anymore and you simplify from there

OpenStudy (anonymous):

yep yep yep right on

OpenStudy (anonymous):

would i subtract 18 and 12?

OpenStudy (anonymous):

get your -6a to be alone.

OpenStudy (anonymous):

you can do that by either subtracting 18 from both sides or by adding 6a and subtracting 12 from both sides.

OpenStudy (anonymous):

18 - 6a = 12 18 - 18 - 6a = 12 - 18 6a = -6 a = 1

OpenStudy (anonymous):

yep good job :) One thing though; you forgot to keep the negative on your -6a. so -6a = -6 a = -6/-6 a = 1 so you were right just forgot to write the -

OpenStudy (anonymous):

ಥ_ಥ so simple, yet so frustrating

OpenStudy (anonymous):

s'all good we're making progress.

OpenStudy (anonymous):

ive learned more on here in the last 10 hours, than i have in algebra since the beginning of the schoool year

OpenStudy (anonymous):

18-6a=12 (subtract 18 from both sides) -6a= -6 (divide through by -6) a= 1